OpenStudy (anonymous):

How many different committees can be formed from 11 teachers and 41 students if the committee consists of 4 teachers and 3 students? In how many ways can the committee of 7 members be selected?

OpenStudy (anonymous):

how do I solve for 11C4? and 41C3?

OpenStudy (anonymous):

???

OpenStudy (kropot72):

The number of combinations of teachers selected 4 at a time is 11C4. The number of combinations of students selected 3 at a time is 41C3 The number of ways of selecting the committee of 7 is given by: \[\frac{11!}{4!7!}\times \frac{41!}{3!38!}=\frac{11\times 10\times 9\times 8}{4\times 3\times 2\times 1}\times\frac{41\times40\times 39}{3\times 2\times 1}=you\ can\ calculate\]

OpenStudy (agent0smith):

\[\Large nCr = \frac{ n! }{ r! (n-r)! }\] so \[\Large 11C4 = \frac{ 11! }{ 4! (11-4)! }\]